Circular No. 8513

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005bm AND 2005bn M. SubbaRao, Department of Astronomy and Astrophysics, University

of Chicago, on behalf of the SDSS Collaboration, reports the discovery of two supernovae in spectra taken as described on IAUC 8302 and 8359. SN 2005bm was detected on Apr. 3.48 UT and is located at α = 15^{h}20^{m}45^{s}.10, δ = +36^{o}48′42′′.5 (equinox 2000.0), which is coincident with the center of the host galaxy whose redshift is z = 0.103; the supernova is not present in an SDSS image of the galaxy taken on 2003 May 1 (limiting magnitudes g = 23.3 and r = 23.1). SN 2005bm appears to be a type-Ia supernova with an approximate age of 4 ± 5 days after maximum light, and its estimated magnitude is r = 19.5 ± 0.2 (the host galaxy has apparent mag r = 17.0 from an image taken prior to the supernova event). SN 2005bn was detected on Apr. 7.36 at α = 12^{h}03^{m}23^{s}.91, δ = +35^{o}19′33′′.0 (also coincident with the center of the host galaxy, which itself has r = 17.9 and z = 0.028). SN 2005bn is not present in an SDSS image of the galaxy taken on 2004 Apr. 13 (same limiting magnitudes as above). The spectrum shows SN 2005bn to be a type-II supernova with an approximate age of 19 ± 5 days after maximum light, and its estimated apparent mag was then also r = 19.5. All supernovae detected as part of this program can be found at

the website http://cheops1.uchicago.edu/pub/, which provides access

to the spectra and finding charts.

SUPERNOVAE 2005bh AND 2005bj N. Morrell, M. Hamuy, G. Folatelli, and C. Contreras, Carnegie Super-

nova Project, report that spectroscopic observations (range 380–930 nm) of SN 2005bh (cf. IAUC 8509) and SN 2005bj (cf. IAUC 8511) were obtained on Apr. 12.14 and 12.35 UT, respectively, with the Las Campanas 2.5-m du Pont telescope (+ WFCCD spectrograph). The spectrum of SN 2005bh is that of a type-Ic supernova and strongly resembles that of SN 1987M at 10 days after maximum light. SN 2005bj is probably also a type-Ic object, quite similar to SN 1994I at 10–12 days after maximum, except for the Si II 635.5-nm line, which is much stronger in the new event.

SUPERNOVAE 2004dj AND 2004et Magnitudes obtained by G. Szabó, University of Szeged, on Feb. 28 at

Konkoly Observatory: SN 2004dj, B = 16.51, V = 15.71, R_{c} = 14.73, I_{c} = 14.30. SN 2004et, B = 17.55, V = 15.97, R_{c} = 14.87, I_{c} = 14.37.

2005 April 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green